You are evaluating a project that costs $77,000 today. The project has an inflow of $164,000 in one year and an outflow of $67,000 in two years.

What are the IRRs for the project? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)

IRR
  Smallest %
  Largest %

  

What discount rate results in the maximum NPV for this project? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  

  IRR %
Answer #1

IRR using the below formula in excel or financial calculator:

1 -77000
2 164000
3 -67000
57.87% IRR(1:3)
1 -77000
2 164000
3 -67000
-44.88% IRR(1:3, -0.99)

Smallest : -44.88%

Largest : 57.87%

b.

Cash inflows * (1+r) 3 = Cash outflows * (1+r) 2

164000 *(1+r) 3 = 144000 * (1+r) 2

1 +r = 0.87804

r = -12.195 %

Note: I have solved part b as per my vest understanding since it says NPV discount but has asked IRR.